# WoW: A Window-to-Window Incremental Index for Range-Filtering Approximate Nearest Neighbor Search

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: the ACM SIGMOD/PODS International Conference on Management of Data; May 31–June 5, 2026; Bengaluru, Indiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Database query processingccs: Information systems Nearest-neighbor search
@@ 1. Introduction
Nearest neighbor search (NNS) (Roussopoulos et al., 1995) is widely used as a fundamental process in various applications. On e-commerce platforms, people use an image to search for similar commodities (Li et al., 2018; Wang et al., 2021b; Wei et al., 2020; Pan et al., 2024). In the retrieval-augmented generation (RAG) pipeline (Gao et al., 2024; Wang et al., 2024), people retrieve paragraphs relevant to a question and provide them as context for large language models. These functions can be realized by encoding the original data into vectors with modern embedding models (Muennighoff et al., 2023; Radford et al., 2021; Bordes et al., 2013; Pennington et al., 2014) and persisting with a vector database management system (VDBMS) (Guo et al., 2022; Pan et al., 2024). With a query also embedded into the same vector space, VDBMS uses $k$-NNS to find similar results in the database. However, finding the exact $k$ nearest neighbors is time-consuming because the time complexity is $O(nd)$, where $n$ is the number of vectors and $d$ is the dimension of vectors (Gao and Long, 2023; Chen et al., 2023). On real-world datasets, the query latency can be even unacceptable for online scenarios, as $n$ varies from millions to billions while $d$ is in the order of hundreds or thousands.
@@
For instance, instead of searching for similar commodities solely by an image, we can provide an additional price range to retrieve affordable ones. Another example resides in a medical question answering system, a typical RAG application (Alkhalaf et al., 2024). When a query comes with “What symptoms are common for people with hypertension aged from 50 to 60?”, the retriever may generate an RFANNS query to find medical records about “hypertension” tagged with ages of patients between 50 and 60.
@@
Figure 1. Example of (a) an oracle proximity graph and (b) a size-4 window graph. Dataset is on the left.
[…]
Example
@@ 3.1. WBT and Hierarchical Window Graphs
Figure 2. Index structure of WoW
[…]
Index structure of WoW
@@ 13    $W_{b}^{l}\leftarrow$ window of size $2o^{l}$ halved by $b$;
Figure 3. Single insertion of WoW. (a) The windows of $v_{73}$ in different layers, and insertion triggers the self-balancing of WBT. (b) The procedure of acquiring candidates for Layer 0 to Layer top, taking Layer 1 for instance. (c) The procedure of selecting neighbors from the candidates and connecting the new vertex into existing graphs, taking Layer 1 for instance.
[…]
WoW insertion
@@ 4.1. Experiment Setting
In our experiments, each query workload contains 1,000 range-filtering queries. We estimate the hardness of a query workload by Local Intrinsic Dimensionality (LID) (Li et al., 2020; Amsaleg et al., 2015; Wang et al., 2025; Aumüller and Ceccarello, 2021), as defined below:
@@ 4.2. RQ1: Index Construction Efficiency
Figure 4. QPS-Recall@10 curves. For each dataset, five workloads with varying range fractions are shown, including the mixed range fraction workloads. Dotted line ($\cdot\cdot\cdot$) stands for static index, dashed line (---) for post-incremental, dash-dotted line (-$\cdot$-) for ordered incremental, and solid line (—) for unordered incremental. Curves below 85% recall are omitted.
[…]
QPS-Recall@10 curves.
@@ 4.3. RQ2: RFANNS Query Performance
    Figure 5. DC-Recall@10 curves compared to oracle HNSW for mixed range fraction. Down and to the right is better.
[…]
    DC-Recall@10 curves compared to oracle HNSW
@@
    Figure 6. Layer range footprints of WoW and w/o early-stop on a single query of Gist under the range fraction $f=2^{-4}$. The results of the first 400 hops are shown. Dark blue bars represent the highest visited layer $l_{max}$ and light ones are the lowest $l_{min}$. The gap between two bars is the exploring depth. The red curve is the trend of the median as $\frac{1}{2}(l_{\min}+l_{\max})$.
[…]
    Layer range footprints
@@
Figure 7. Effectiveness of landing layer selection. The layers selected by Algorithm 3 are given on the top right corner. Some curves are absent for failing to achieve 85% recall.
[…]
Effectiveness of landing layer selection
@@ 4.4. RQ3: Detailed Analysis
Figure 8. QPS-Recall@10 under workloads with different query correlations on the Sift dataset
[…]
QPS-Recall@10 under workloads with different query correlations on the Sift dataset
Figure 9. DC and Recall@10 for different LID@10 (marked in the middle of the bars)
[…]
DC and Recall@10 for different LID@10
@@
Figure 10. QPS-Recall@$k$ ($k=1,50,100$) on the mixed range fraction workloads of the Sift and Gist datasets
[…]
QPS-Recall@$k$ ($k=1,50,100$) on the mixed range fraction workloads of the Sift and Gist datasets
Figure 11. Parameter sensitivity on all datasets
[…]
Parameter sensitivity on all datasets
Figure 12. QPS-Recall@10 with varying numbers of unique values. Range filters are generated randomly with mixed selectivity.
[…]
QPS-Recall@10 with varying numbers of unique values.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No. 62272219). We appreciate Mr. Shuo Shen for his contributions to deploying baseline methods.
@@ 3 $i\leftarrow\mathrm{GetRank}(x_{u})$, $j\leftarrow\mathrm{GetRank}(y_{l})$;
5 procedure $\mathrm{GetRank}(a)$

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: 10.1145/3769843conference: the ACM SIGMOD/PODS International Conference on Management of Data; May 31–June 5, 2026; Bengaluru, Indiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Database query processingccs: Information systems Nearest-neighbor search
@@ 1. Introduction
Nearest neighbor search (NNS) (Roussopoulos et al., 1995) is widely used as a fundamental process in various applications. On e-commerce platforms, people use an image to search for similar commodities (Li et al., 2018; Wang et al., 2021b; Wei et al., 2020; Pan et al., 2024; Wang et al., 2025a). In the retrieval-augmented generation (RAG) pipeline (Gao et al., 2024; Wang et al., 2024), people retrieve paragraphs relevant to a question and provide them as context for large language models. These functions can be realized by encoding the original data into vectors with modern embedding models (Muennighoff et al., 2023; Radford et al., 2021; Bordes et al., 2013; Pennington et al., 2014) and persisting with a vector database management system (VDBMS) (Guo et al., 2022; Pan et al., 2024). With a query also embedded into the same vector space, VDBMS uses $k$-NNS to find similar results in the database. However, finding the exact $k$ nearest neighbors is time-consuming because the time complexity is $O(nd)$, where $n$ is the number of vectors and $d$ is the dimension of vectors (Gao and Long, 2023; Chen et al., 2023). On real-world datasets, the query latency can be even unacceptable for online scenarios, as $n$ varies from millions to billions while $d$ is in the order of hundreds or thousands.
@@
For instance, instead of searching for similar commodities solely by an image, we can provide an additional price range to retrieve affordable ones. Another example resides in a medical question answering system, a typical RAG application (Alkhalaf et al., 2024; Sabbir M. Rashid, 2025). When a query comes with “What symptoms are common for people with hypertension aged from 50 to 60?”, the retriever may generate an RFANNS query to find medical records about “hypertension” tagged with ages of patients between 50 and 60.
@@
Figure 1. Example of (a) an oracle proximity graph and (b) a size-4 window graph. Dataset is on the left.:
@@ 3.1. WBT and Hierarchical Window Graphs
Figure 2. Index structure of WoW:
@@ 13    $W_{b}^{l}\leftarrow$ window of size $2o^{l}$ halved by $b$;
Figure 3. Single insertion of WoW. (a) The windows of $v_{73}$ in different layers, and insertion triggers the self-balancing of WBT. (b) The procedure of acquiring candidates for Layer 0 to Layer top, taking Layer 1 for instance. (c) The procedure of selecting neighbors from the candidates and connecting the new vertex into existing graphs, taking Layer 1 for instance. :
@@ 4.1. Experiment Setting
In our experiments, each query workload contains 1,000 range-filtering queries. We estimate the hardness of a query workload by Local Intrinsic Dimensionality (LID) (Li et al., 2020; Amsaleg et al., 2015; Wang et al., 2025b; Aumüller and Ceccarello, 2021), as defined below:
@@ 4.2. RQ1: Index Construction Efficiency
Figure 4. QPS-Recall@10 curves. For each dataset, five workloads with varying range fractions are shown, including the mixed range fraction workloads. Dotted line ($\cdot\cdot\cdot$) stands for static index, dashed line (---) for post-incremental, dash-dotted line (-$\cdot$-) for ordered incremental, and solid line (—) for unordered incremental. Curves below 85% recall are omitted.:
@@ 4.3. RQ2: RFANNS Query Performance
    Figure 5. DC-Recall@10 curves compared to oracle HNSW for mixed range fraction. Down and to the right is better.:
@@
    Figure 6. Layer range footprints of WoW and w/o early-stop on a single query of Gist under the range fraction $f=2^{-4}$. The results of the first 400 hops are shown. Dark blue bars represent the highest visited layer $l_{max}$ and light ones are the lowest $l_{min}$. The gap between two bars is the exploring depth. The red curve is the trend of the median as $\frac{1}{2}(l_{\min}+l_{\max})$.:
@@
Figure 7. Effectiveness of landing layer selection. The layers selected by Algorithm 3 are given on the top right corner. Some curves are absent for failing to achieve 85% recall.:
@@ 4.4. RQ3: Detailed Analysis
Figure 8. QPS-Recall@10 under workloads with different query correlations on the Sift dataset:
Figure 9. DC and Recall@10 for different LID@10 (marked in the middle of the bars):
@@
Figure 10. QPS-Recall@$k$ ($k=1,50,100$) on the mixed range fraction workloads of the Sift and Gist datasets:
Figure 11. Parameter sensitivity on all datasets:
Figure 12. QPS-Recall@10 with varying numbers of unique values. Range filters are generated randomly with mixed selectivity.:
@@ Acknowledgements.
This work was supported by the Noncommunicable Chronic Diseases - National Science and Technology Major Project (Grant No.: 2023ZD0503600/2023ZD0503604). We thank Mr. Shuo Shen for his contributions to deploying baseline methods.
@@ 3 $i\leftarrow\mathrm{GetRank}(x_{u})$, $j\leftarrow\mathrm{GetRank}(y_{l})$;
5
61ex procedure $\mathrm{GetRank}(a)$